OpenStudy (anonymous):

if a cinema took in $1270 in ticket sales for a movie that 220 people attended. child tickets are $4.50 and adults ate $6.50, how many children attended?

OpenStudy (sandra):

there are two unknowns, the number of child tickets sold and the number of adult tickets sold

OpenStudy (sandra):

so you know you need two equations to solve it

OpenStudy (sandra):

1. c + a = 220 2. 4.5c + 6.5a = 1270 3. Now you want to take the value of a in the first equation, and substitute it into the seocnd

OpenStudy (sandra):

so a = 220 - c

OpenStudy (sandra):

4. 4.5c + 6.5* (220 -c) = 1270

OpenStudy (sandra):

5. 4.5c - 6.5c + 1430 = 1270

OpenStudy (sandra):

6. now get all of the c's on one side, and the plane numbers on the other

OpenStudy (sandra):

so 1430 - 1270 = 6.5c - 4.5c 160 = 2c c = 80